Present value. Prestigious University is offering a new adr Prest gous f they pay the first years tuition. The university will pay an annual rate of return of 5 5% on the deposited tution, and a full refund will be all the child chooses another university. The tuition is expected to be $14,000 a year at Prestigious 20 years from now. What would parents pay today if they just gave birth to a new baby and the child will attend college in 20 years? How much is the required payment to secure admission for their child if the interest rate falls to 490 sion and tuition payment plan for all alumni. On the birth of a child, parents can guarantee admission to
Answer #1

Tuition fee in 20 years from now = $14,000

Interest Rate = 5.5%

To calculate what parents need to pay today,

PV = FV/(1+r)t

PV = 14000/(1.055)20

PV = $4,798.20

So parents need to pay $4,798.20 today for tuition fee after 20 years.

If interest rate falls to 4%

PV = 14000/(1.04)20

PV = $6,389.42

So, if interest rate falls to 4% parents need to pay $6,389.42 today for tuition fee after 20 years.